### v1

# A Morphology-Based Investigation of Positional Encodings

Abstract
How does the importance of positional encoding in pre-trained language models (PLMs) vary across languages with different morphological complexity? In this paper, we offer the first study addressing this question, encompassing 23 morphologically diverse languages and 5 different downstream tasks. We choose two categories of tasks: syntactic tasks (part-of-speech tagging, named entity recognition, dependency parsing) and semantic tasks (natural language inference, paraphrasing). We consider language-specific BERT models trained on monolingual corpus for our investigation. The main experiment consists of nullifying the effect of positional encoding during fine-tuning and investigating its impact across various tasks and languages. We perform additional experiments on a subset of tasks and languages, where we deliberately distort the original word order of a sentence. Our findings demonstrate that the significance of positional encoding diminishes as the morphological complexity of a language increases. Across all experiments, we observe clustering of languages according to their morphological typology - with analytic languages at one end and synthetic languages at the opposite end.
Keywords: Pre-trained language models (PLMs), Positional Encoding, Morphosyntax, Linguistic Diversity
[…]
A Morphology-Based Investigation of Positional Encodings
Poulami ${}^{\clubsuit}$, Shikhar ${}^{\spadesuit}$, Raj ${}^{\diamondsuit}$, Pushpak ${}^{\clubsuit}$
${}^{\clubsuit}$omputer Science and Engineering, IIT Bombay, India
${}^{\spadesuit}$oogle Research, India ${}^{\diamondsuit}$, Japan
{poulami, pb}@cse.iitb.ac.in
Abstract content
1.   Introduction
Transformers Vaswani et al. (2017) have ushered in a new era in the field of natural language processing, facilitating the development of highly versatile language models (LMs) that have achieved ground-breaking results across a wide spectrum of language processing tasks such as natural language inference Liu et al. (2019a), text classification Raffel et al. (2020a), named entity recognition Liu et al. (2019a), and part-of-speech tagging Martin et al. (2020). This has led to the emergence of various pre-trained language models Lan et al. (2019); Xue et al. (2020); Brown et al. (2020b). The training of these models on vast quantities of raw text equips them with the ability to accurately capture the inherent linguistic features of the language Tenney et al. (2019). The general paradigm is to further fine-tune a pre-trained model on task-specific or domain-specific data for any downstream task.
However, existing models often overlook the diverse fine-grained and broader domains of linguistics. This is particularly evident since there has been very few instances where morphology of a language has been taken into consideration while designing a language model. For instance, Nzeyimana and Rubungo (2022) proposes a morphology-aware language model build on a two-tier transformer architecture for low-resource morphologically rich Kinyarwanda language. Park et al. (2021) study the impact of morphological complexity on LSTM based language models. As language models are being developed for more and more languages which significantly differ in morphological typology, it is imperative to incorporate linguistic knowledge and construct language models that are sensitive to these linguistic nuances. This serves as the motivation behind our work in this paper.
Figure 1: Methodology Overview
Figure 2: The figure illustrates the effect of word order on semantics for two languages: English (left) and Sanskrit (right). English is a morphologically poor language with SVO word order whereas Sanskrit is a morphologically rich language with no dominant word order(NODOM). Distorting the word order completely alters the meaning for English. However, for Sanskrit the meaning remains intact. More details in Section 4.1.
The vast diversity of morphological systems exhibited in languages worldwide constitutes an intriguing dimension of human language. The central focus of our paper revolves around investigating the relationship between morphology and word order by leveraging the concept of positional encoding. Our contributions are:
  • We perform the first study about the varying importance of positional encoding across languages with different morphological complexity.
  • We are the first to show that the impact of positional encodings diminishes as the morphological complexity of a language increases.
  • We conduct exhaustive experiments covering 23 different languages across 8 language families and 5 diverse natural language processing tasks.
  • Across all results, we observe clustering of languages according to their morphological typology with analytic languages at one end and synthetic languages at the opposite end.
2.   Related work
In this section, we provide a comprehensive overview of pre-trained language models, positional encodings and how they relate to linguistics.
Pre-trained language models: Pre-trained language models such as BERT(Devlin et al., 2018), RoBERTa(Liu et al., 2019b), GPT-2(Radford et al., 2019), T5(Raffel et al., 2020b), GPT-3(Brown et al., 2020a) have led to significant advancements in the field of natural language processing. However, the enormous computational cost incurred during their training is a major challenge in the development of pre-trained language models. Acquiring a deeper understanding of how various components of a pre-trained language model function in different languages can provide valuable insights regarding their necessity (or redundancy) across languages, helping in the efficient reduction of computational costs.
Positional Encodings: Positional encoding is a crucial component in transformer-based language models. It is used to explicitly add positional information and capture word order within a sequence. There exist several variants of positional embeddings such as absolute positional embeddings Vaswani et al. (2017); Radford et al. (2019), and relative positional embeddings Shaw et al. (2018); Raffel et al. (2020b); Devlin et al. (2018). They were originally introduced for English, a morphologically poor language. Thus, it is uncertain whether their effectiveness can be generalized across languages with varying morphological complexity.
Probing pre-trained language models for linguistic knowledge: As PLMs are being developed for more and more languages that significantly differ in linguistic complexity, it is imperative to incorporate linguistic knowledge and construct language models that are sensitive to different linguistic nuances. Recently, there has been a surge in the study of linguistic knowledge encoded by pre-trained language models, focusing different aspects of linguistics such as syntax (Goldberg, 2019; Jawahar et al., 2019) and semantics (Ethayarajh, 2019). There has also been several studies to understand how morphology is encoded in PLMs. Sadeniemi et al. (2008) investigate the morphological, morpho-syntactic and syntactic complexities of European Union languages by using a compression based metric motivated by Kolmogorov complexity. Gerz et al. (2018) investigate the connection between language modeling and linguistic typology across 50 different languages. However their work do not include transformer-based language models and encompass a limited set of tasks. Edmiston (2020) probe BERT models of five European languages for morphological information. Otmakhova et al. (2022) analyze how different layers of the Transformer encode morphology and syntax. But they focus their study only on three languages and devise a simple classification task such as identification of word order corruption to understand the correlation between word order and morphology in BERT.
Infusing linguistic knowledge in language models: Over the past few years, there has been significant efforts to infuse linguistic knowledge into pre-trained language models. Chaudhary et al. (2018) adapt word embeddings by leveraging morphological and phonological information. Nzeyimana and Niyongabo Rubungo (2022) employed feature-based pre-training to design a morphology-aware language model for a low-resource African language, Kinyarwanda. Chakrabarty et al. (2022) propose a encoder-decoder pre-trained model enriched with linguistic features such as lemma, part-of-speech and dependency labels. All these works show that incorporating linguistic knowledge improves the performance of pre-trained models in downstream tasks. However, they don’t shed light on how the relevance of different components of a transformer model can vary across languages.
3.   The Relationship between Morphology and Word Order
In this section, we investigate the relationship between morphology and word order as outlined in different linguistic theories.
Linguistic theories : Morphological case markings serve a similar function as word order(Sapir, 1921; Blake, 2001).Several theories suggest that the presence of morphological case is necessary for free word order in a language (Comrie, 1989; Haspelmath, 1999). Either morphological case or structural position facilitates the unambiguous determination of grammatical role of the constituents of a sentence. The existence of morphological case reduces the need for fixed structural position in determination of the grammatical function of a word or phrase, allowing for variable word order. However, if morphological case is absent, fixed placement of words (and phrases) is necessary, exhibiting a fixed or rigid word order. In this work, we align our empirical study in accordance with the above theories that hints at the existence of a correlation between morphology and word order. Specifically, morphologically rich languages which tend to exhibit higher word-order flexibility as compared to morphologically poor languages [^1] The theory regarding morphology and word order is a linguistically complex topic. Concurrent theories in the literature propose that there is no correlation between morphological complexity and word order. Müller (2002) demonstrates phenomena like scrambling and topic shift, where the change in word order does not necessarily require a high level of morphological complexity. However, this is beyond the scope of our study..
3.1.   Spectrum of Morphological Complexity
Through the lens of morphological typology Haspelmath and Sims (2013), we can categorize and cluster languages by studying their inherent morphological structures. At one extreme, we find languages such as Chinese and Vietnamese, which fall into the category of analytic languages and are morphologically impoverished. In these languages, it is essential for words to maintain fixed positions in order to accurately convey grammatical relationships, resulting in a strict and invariant word order. On the other extreme, we find synthetic languages such as Sanskrit and Finnish, known for their rich morphology, where it’s possible to rearrange the word order within a sentence without changing its meaning, as illustrated in Figure 2. However, most languages fall between these two extremes. Synthetic languages can be categorized into two main types. Agglutinative languages like Hungarian and Turkish tend to stick together multiple morphemes while fusional languages fuse several morphemes to express various grammatical features.
4.   Methodology
4.1.   Morphology-based Investigation
In Transformer-based Vaswani et al. (2017) language models such as BERT Devlin et al. (2018), a given input sequence $S$ consisting of tokens $w_{1},w_{2},...w_{n}$ is fed to an embedding layer to generate the corresponding word embeddings $\bm{e}_{1},\bm{e}_{2},...,\bm{e}_{n}$. This is followed by multiple self-attention layers. However, the attention mechanism is completely invariant to the order of tokens which is crucial for capturing semantics in languages such as English. Figure 2 illustrates that how the same set of tokens in English can have completely different meaning given their different ordering. Hence, positional embeddings are added to the word embeddings to effectively capture the ordering of tokens. For each token $w_{i}$, a real-valued vector $\bm{p}_{i}\in\mathbb{R}^{d}$ is added to its token embedding $\bm{e}_{i}$. Thus, a position-aware token embedding $\bm{e}^{\prime}_{i}=\bm{e}_{i}+\bm{p}_{i}$ is fed to the self-attention stack. Vaswani et al. (2017) proposes the use of sine and cosine functions of different frequencies to encode positional information.
$PE_{(pos,2i)}=sin\left(\frac{pos}{10000^{\frac{2i}{d_{model}}}}\right)$
$PE_{(pos,2i+1)}=cos\left(\frac{pos}{10000^{\frac{2i}{d_{model}}}}\right)$
where $pos$ represents the position of a token in a sequence and $i$ denotes the $i^{th}$ dimension of the embedding vector. To study the impact of positional embeddings on Transformer-based LMs for languages with different morphological richness, we propose the following two strategies.
  • Nullifying Positional Encoding: In this strategy, we set positional embedding in the language model to null, i.e., $\bm{p}_{i}=\bm{0}$. Thus, making model invariant to the token word of the input sequence. We hypothesize that for morphological rich languages such as Sanskrit this would not have much impact on the downstream performance. For instance, in Figure 2, we demonstrate that for Sanskrit, the meaning of the sentence remains consistent even when the token order is shuffled. However, this does not hold for morphologically poor languages such as English.
  • Permuting Word Order: Here, we distort the word order by shuffling a sentence at the level of uni-gram, bi-gram and tri-gram. The primary objective is to introduce varying degrees of distortion and examine the resulting loss in information across morphologically diverse languages. The proposed permutation strategy can be applied to study the robustness of the language models to the token order of the input.
5.   Experimental Setup
To ensure the generalizability of our findings, we choose to perform a comprehensive study spanning different languages and tasks. In this section, we outline the principles for selecting diverse tasks and languages and provide details on the datasets and models used in evaluation.
5.1.   Tasks
The main objective of our research is to analyze the effect of positional encoding in languages with varying morphological complexity. As discussed in Section 2, positional encoding encodes word order and thereby the syntactic relationships inherent in a language. Since our work deals with the interplay of morphology and syntax in pre-trained language models, we consider two sets of tasks:
  • Syntactic tasks - Part-of-speech tagging, Named Entity Recognition, Dependency Parsing
  • Semantic tasks - Natural Language Inference(NLI), Paraphrasing
This choice of tasks enables us to understand the function of morphology and word order in encoding syntactic and semantic role information across various languages. Moreover, text classification tasks such as natural language inference and paraphrasing operate at the sentence level. On the other hand, part-of-speech tagging and named entity recognition are sequence labeling tasks that assess the model’s competence at a more granular word level. Additional factors considered in task selection include:
  • [itemsep=2pt,parsep=2pt,partopsep=2pt,leftmargin=*,topsep=2pt]
  • Multilinguality: We select tasks that span a diverse set of languages and language families to establish the broad applicability of our results.
  • Sufficient monolingual training data: We consider tasks with sufficient monolingual training data. As we remove positional encoding at the time of fine-tuning, the training data needs to be substantial for the impact of positional encoding removal to be profound.
5.2.   Languages
The choice of our languages is determined by the fulfillment of the following criteria: (1) availability of monolingual BERT-base model on HuggingFace Hub (Wolf et al., 2019), (2) availability of sufficient monolingual training data across different tasks and (3) typological diversity. We aim to cover as many languages and language families as possible. Overall, we cover 23 languages distributed across 8 language families and one language isolate. In Table 1, we present an outline of these languages
Language(ISO code) Language family Hugging Face Model id
Arabic (ar) Afro-Asiatic aubmindlab/bert-base-arabertv02
Basque (eu) Basque (Language isolate) orai-nlp/ElhBERTeu
Bulgarian (bg) Indo-European: Slavic usmiva/bert-web-bg
Chinese (zh) Sino-Tibetan bert-base-chinese
English (en) Indo-European: Germanic bert-base-cased
Finnish (fi) Uralic TurkuNLP/bert-base-finnish-cased-v1
French (fr) Indo-European: Romance dbmdz/bert-base-french-europeana-cased
German (de) Indo-European: Germanic dbmdz/bert-base-german-cased
Greek (el) Indo-European: Greek nlpaueb/bert-base-greek-uncased-v1
Hebrew (he) Afro-Asiatic onlplab/alephbert-base
Hindi (hi) Indo-European: Indo-Aryan l3cube-pune/hindi-bert-scratch
Hungarian (hu) Uralic SZTAKI-HLT/hubert-base-cc
Indonesian (id) Austronesian indolem/indobert-base-uncased
Italian (it) Indo-European: Romance dbmdz/bert-base-italian-cased
Korean (ko) Koreanic kykim/bert-kor-base
Malay (ms) Austronesian malay-huggingface/bert-base-bahasa-cased
Persian (fa) Indo-European: Iranian HooshvareLab/bert-base-parsbert-uncased
Portuguese (pt) Indo-European: Romance neuralmind/bert-base-portuguese-cased
Russian (ru) Indo-European: Slavic DeepPavlov/rubert-base-cased
Spanish (es) Indo-European: Romance dccuchile/bert-base-spanish-wwm-cased
Swahili (sw) Niger-Congo flax-community/bert-base-uncased-swahili
Turkish (tr) Turkic dbmdz/bert-base-turkish-cased
Vietnamese (vi) Austro-Asiatic trituenhantaoio/bert-base-vietnamese-uncased
Table 1: Overview of different languages
5.3.   Datasets
In our experiments, we evaluate our models on a variety of tasks. We utilize GLUE Wang et al. (2018) benchmark which comprises of nine language understanding tasks encompassing a wide spectrum of domains. The benchmark includes sentence acceptability judgment on CoLA Warstadt et al. (2018), sentiment analysis on SST-2 Socher et al. (2013), paraphrasing/sentence similarity on MRPC Dolan and Brockett (2005), STS-B Cer et al. (2017), QQP Sharma et al. (2019), and natural language inference on MNLI Williams et al. (2018), QNLI Rajpurkar et al. (2016), and RTE Bentivogli et al. (2009)) datasets.
In order to broaden the scope of our analysis beyond the English language, we incorporate multilingual datasets that span various domains and encompass multiple languages. Our investigation includes tasks from the XTREME benchmark Hu et al. (2020), covering natural language inference (XNLI) Conneau et al. (2018), paraphrasing (PAWS-X) Yang et al. (2019), and structure prediction tasks such as part-of-speech (POS) tagging and named entity recognition (NER). We use the data from the Universal Dependencies v2.12 (Zeman et al., 2023) for the task of dependency parsing. The treebanks used for different languages are specified in Table 3. By incorporating these tasks, we aim to enhance the comprehensiveness and breadth of our research.
5.4.   Evalution Metrics
5.5.   Models
In this section, we describe the baseline and perturbed models used in our experiments.
5.6.   Model Selection
In our research, we explore monolingual pre-trained language models to prevent cross-lingual transfer from influencing our results. Given the availability of monolingual BERT models in various languages, we select BERT as the example PLM for our study. We consider BERT-base model for all languages to ensure that variations in model size and architecture do not influence the results.
5.6.1.   Baseline Models
We used fine-tuned BERT-base models as the baseline.
5.6.2.   Perturbed Models
We conduct experiments on two perturbed models, as outlined in Section 4.1.
  • Models without positional encoding: During the fine-tuning stage, we nullify positional encoding and measure the relative drop in performance with respect to the baseline model to assess how positional encoding impacts different languages. [^2] Training language models from scratch without positional encoding requires significant computational power and financial resources. Hence, due to limited computational resources, we focused on fine-tuning rather than pre-training. We discuss this further in section Limitations.
  • Models with perturbed input: Given a sentence, we compute a list of N-grams and randomly shuffle the entire list to generate the distorted sentence. We ensure that the shuffled sentence is different from the original sentence. We use non-perturbed training and development sets and only perturb the test inputs. We measure the relative drop in performance with respect to the baseline model to observe how word order corruption impacts different languages.
5.7.   Training and evaluation setup
For text classification tasks like natural language inference, paraphrasing and sequence labelling tasks like named-entity recognition and POS tagging, we adhered to the generic pipeline. We implemented a biaffine parser for dependency parsing, by applying a biaffine attention layer directly on the output of BERT as described in Glavaš and Vulić (2021). As outlined in Devlin et al. (2018), we have performed hyper-parameter tuning and fine-tuned BERT on GLUE tasks. For multilingual experiments, hyper-parameter tuning is performed on English validation data as suggested in the XTREME benchmark. However, since our goal is not to achieve the best absolute performance, we avoided conducting extensive hyperparameter tuning. For the reproducibility of our results, we will add the hyperparameter details in the appendix of the paper.
Task Dataset With Positional Encoding Without Positional Encoding Relative decrease
Unigram Bigram Trigram Without
Permutation Permutation Permutation Permutation
Sentence
Acceptability
CoLA 4.4 12.3 16.8 59.1 23.8 59.7
Sentiment
Analysis
SST-2 81.6 86.0 85.1 91.8 86.5 5.8
Paraphrasing / Sentence Similarity MPRC 83.5 84.2 85.4 89.8 84.6 5.8
QQP 75.6 79.1 80.8 87.1 85.8 1.5
STS-B 85.2 87.1 86.6 89.0 86.6 2.7
Natural Language Inference MNLI 68.3 74.8 76.5 83.6 79.7 4.7
MNLI-MM 68.7 74.4 76.6 84.0 79.8 5.0
QNLI 81.3 85.0 86.5 91.0 87.2 4.2
RTE 58.1 61.5 61.8 64.5 62.8 2.6
Table 2: GLUE Results : The evaluation metrics used for reporting the performance of QQP and MRPC tasks are F1 scores, while for the STS-B task, Spearman correlations are used, and accuracy scores are employed for the remaining tasks. The average and standard deviation are reported across 3 trails of the evaluation on the validation set. The relative decrease quantifies the decline in performance when positional encoding is excluded compared to when positional encoding was present. Additionally, we conducted experiments in which we removed positional encoding and perturbed the input to the model. Since the removal of positional encoding results in a bag of words model, we observed no noticeable change upon further distortion.
6.   Results
We conduct preliminary experiments on the GLUE benchmark, designed for English, and subsequently expand our study to encompass multiple languages.
6.1.   Impact of positional encoding
6.1.1.   Results on GLUE tasks
Removing positional encoding leads to a varied decrease in performance across different tasks, as evident in Table 2.
Sentence/Grammatical Acceptability Tasks: Positional encoding helps the model understand the hierarchical structure and dependencies between words, which is essential for determining the grammaticality of a sentence. As a result, in case of CoLA task, when positional encoding is removed, the model struggles to identify grammatically acceptable sentences, leading to a notable decline of 59.7% in performance.
Paraphrasing and sentence similarity Tasks: Models can effectively capture the similarity or relatedness between sentences by focusing on common signals present across sentences. These tasks primarily require understanding the underlying semantic meaning and contextual similarities between sentences rather than the syntactic structure.As a result, when positional encoding is removed, the relative decrease in performance is considerably smaller (5.8% for MRPC, 2.7% for STS-B, and 1.5% for QQP). This indicates that while positional encoding does provide some benefit in capturing the positional information within sentences, it is not very crucial for these tasks.
Natural language inference Tasks: The removal of positional encoding leads to a relative decrease of 4.7% for MNLI, 5% for the mismatched version of MNLI, 4.2% for QNLI, and 2.6% for RTE. The decrease in performance is still relatively modest for these tasks. This highlights that positional encoding does not play a significant role in understanding the logical inference and entailment relationships between sentence pairs.
Even in the absence of positional encoding, the bag of words model gives considerably good performance for paraphrasing and natural language inference tasks.The results on GLUE benchmark serve as a driving force behind our investigation, where we aim to further test our hypothesis across morphologically diverse languages.
6.1.2.   Results on Syntactic Tasks
Figures 3,4 and 5 demonstrate the effect of removing positional encoding in part-of-speech tagging, named entity recognition and dependency parsing tasks.
  • The relative drop in performance due to removal of positional encoding decreases with increasing morphological complexity.
  • Analytic languages like Chinese and Vietnamese, characterized by minimal or no morphology, exhibit the most significant decrease in performance when positional encoding is removed and cluster at one end. Moderately analytic languages like English and French follow.
  • At the opposite end of the spectrum, we find synthetic languages such as Hungarian, Finnish, and Turkish, known for their rich morphological systems. In these languages, the function of morphology in encoding grammatical roles surpasses that of word order, resulting in a considerably smaller decrease in performance when positional encoding is eliminated.
It is important to note that the methods employed for data annotation in the XTREME benchmark for POS tagging and NER datasets are different. While the POS tagging dataset was human annotated, the NER dataset was created through automatic annotation using weak supervision. Despite these disparities, the findings of the POS tagging and NER experiments are similar.
6.1.3.   Results on Semantic Tasks
Natural language inference and paraphrasing tasks are not inherently focused on the study of syntax or grammatical structure. Instead, these tasks primarily involve understanding the semantic relationships and meaning between sentences or phrases. While syntax can play a role in these tasks to some extent, the main emphasis is on capturing the semantic equivalence or entailment between sentences and generating accurate paraphrases. Syntax, including word order and grammatical rules, can influence the overall coherence and clarity of the expressions, but it is not the primary focus of these tasks. We employ the XNLI and PAWS-X datasets from the XTREME benchmark for natural language inference and paraphrasing tasks.The results depicted in Figures 8 and 9 illustrate the consequences of eliminating positional encoding in tasks related to natural language inference and paraphrasing. We notice a consistent pattern emerge in the graphs where morphologically poor languages are notably affected by the absence of positional encoding, while the impact is comparatively less for morphologically rich languages. However, in contrast to syntactic tasks, the variability in impact across different languages is less pronounced for semantic tasks.
6.2.   Impact of permutation:
6.2.1.   Results on GLUE tasks
In addition to studying the effect of positional encoding, we also conducted experiments to examine the impact of permutation on various GLUE tasks.
  • [itemsep=2pt,parsep=2pt,partopsep=2pt,leftmargin=*,topsep=2pt]
  • Unigram permutation causes the most significant performance drop. However, as we increase the ngram order, which involves shuffling chunk of words instead of individual words, the decrease in performance is significantly less. This indicates that shuffling at higher ngram levels add less distortion and preserve the integrity of word order to a greater extent.
  • The results also imply that lower order ngrams capture vocabulary match and is completely ignorant of word order whereas higher order ngrams capture word order and other dependencies present in a sentence.
Figure 3: Effect of Positional Encoding on POS Tagging.
Figure 4: Effect of Positional Encoding on NER task.
Figure 5: Effect of Positional Encoding on Dependency Parsing.
Figure 6: Effect of Permutation on NER task.
Figure 7: Effect of Permutation on POS tagging task.
Figure 8: Effect of Positional Encoding on XNLI.
Figure 9: Effect of Positional Encoding on PAWS-X.
6.2.2.   Results on Syntactic tasks
We empirically measure the information context preserved in word structure by studying the degradation in performance of downstream tasks when the word order is distorted. Figures 6 and 7 illustrate the impact of n-gram word order shuffling on named entity recognition and part-of-speech tagging tasks.
  • Figures 6 and 7 indicate that the distortion of word order results in minimal information loss for languages with rich morphological systems as compared to languages with impoverished morphology.
  • For morphologically poor languages, there is a drastic drop in performance since shuffling the word order either alters the meaning of a sentence or produces completely incoherent statements for morphologically poor languages. Since morphologically poor languages majorly depend on word order, removal of positional encoding would deprive the model of its knowledge of syntax,leaving the model ungrammatical.
  • In both the figures 6 and 7 morphologically poor languages cluster at one end, while morphologically rich languages cluster at the opposite end. As we move along the graph, there is a gradual increase in the morphological complexity of languages. This leads to a higher degree of flexibility in their word order and a stronger dependence on morphology. Consequently, the influence of word order permutation diminishes as languages become more morphologically complex.
Language Treebank
Chinese (zh) UD_Chinese-GSD
Portuguese (pt) UD_Portuguese-Bosque
Spanish (es) UD_Spanish-GSD
English (en) UD_English-GUM
French (fr) UD_French-GSD
Italian (it) UD_Italian-ISDT
Russian (ru) UD_Russian-Taiga
German (de) UD_German-GSD
Basque (eu) UD_Basque-BDT
Finnish (fi) UD_Finnish-FTB
Turkish (tr) UD_Turkish-Penn
Table 3: Details of treebanks of different languages
7.   Conclusion
In this work, we demonstrate the interplay between positional encoding and morphology for various morphologically diverse languages. We present the first study regarding the varying impact of positional encoding across languages with varying morphological complexity. Our results reveal that the importance of positional encoding diminishes as the morphological complexity of a language increases. In all our experiments, we notice a clustering of languages based on their morphological typology, with analytic languages on one end and synthetic languages on the other end. By understanding how positional encoding and morphology affect model performance, developers can make more informed decisions about which techniques to use for specific tasks and create more robust models.
Ethics Statement
All the experiments are conducted on openly available datasets and benchmarks with no ethical consideration.
Limitations
Our goal was to incorporate multiple languages to ensure the generalizability of our findings. However, the process of training language models from scratch (without positional encoding) for a large number of languages requires significant computational power and financial resources. We agree that pre-training without positional encoding would be a more holistic approach. However, due to limited computational resources, our primary focus was directed towards fine-tuning. However, we firmly believe that removing positional encoding in the pre-training phase would yield more pronounced results.
8.   Bibliographical References
[…]
  • Aho and Ullman (1972) Alfred V. Aho and Jeffrey D. Ullman. 1972. The Theory of Parsing, Translation and Compiling, volume 1. Prentice-Hall, Englewood Cliffs, NJ.
  • American Psychological Association (1983) American Psychological Association. 1983. Publications Manual. American Psychological Association, Washington, DC.
  • Ando and Zhang (2005) Rie Kubota Ando and Tong Zhang. 2005. A framework for learning predictive structures from multiple tasks and unlabeled data. Journal of Machine Learning Research, 6:1817–1853.
  • Andrew and Gao (2007) Galen Andrew and Jianfeng Gao. 2007. Scalable training of $L_{1}$-regularized log-linear models. In Proceedings of the 24th International Conference on Machine Learning, pages 33–40.
  • Bahdanau et al. (2014) Dzmitry Bahdanau, Kyunghyun Cho, and Yoshua Bengio. 2014. Neural machine translation by jointly learning to align and translate. arXiv preprint arXiv:1409.0473.
  • Bentivogli et al. (2009) Luisa Bentivogli, Peter Clark, Ido Dagan, and Danilo Giampiccolo. 2009. The sixth pascal recognizing textual entailment challenge. In Text Analysis Conference.
  • Blake (2001) Barry J Blake. 2001. Case. Cambridge University Press.
  • Brown et al. (2020a) Tom Brown, Benjamin Mann, Nick Ryder, Melanie Subbiah, Jared D Kaplan, Prafulla Dhariwal, Arvind Neelakantan, Pranav Shyam, Girish Sastry, Amanda Askell, et al. 2020a. Language models are few-shot learners. Advances in neural information processing systems, 33:1877–1901.
  • Brown et al. (2020b) Tom B. Brown, Benjamin Mann, Nick Ryder, Melanie Subbiah, Jared Kaplan, Prafulla Dhariwal, Arvind Neelakantan, Pranav Shyam, Girish Sastry, Amanda Askell, Sandhini Agarwal, Ariel Herbert-Voss, Gretchen Krueger, T. J. Henighan, Rewon Child, Aditya Ramesh, Daniel M. Ziegler, Jeff Wu, Clemens Winter, Christopher Hesse, Mark Chen, Eric Sigler, Mateusz Litwin, Scott Gray, Benjamin Chess, Jack Clark, Christopher Berner, Sam McCandlish, Alec Radford, Ilya Sutskever, and Dario Amodei. 2020b. Language models are few-shot learners. ArXiv, abs/2005.14165.
  • BSI (1973a) BSI. 1973a. Natural Fibre Twines, 3rd edition. British Standards Institution, London. BS 2570.
  • BSI (1973b) BSI. 1973b. Natural fibre twines. BS 2570, British Standards Institution, London. 3rd. edn.
  • Castor and Pollux (1992) A. Castor and L. E. Pollux. 1992. The use of user modelling to guide inference and learning. Applied Intelligence, 2(1):37–53.
  • Cer et al. (2017) Daniel Cer, Mona Diab, Eneko Agirre, Iñigo Lopez-Gazpio, and Lucia Specia. 2017. SemEval-2017 task 1: Semantic textual similarity multilingual and crosslingual focused evaluation. In Proceedings of the 11th International Workshop on Semantic Evaluation (SemEval-2017), pages 1–14, Vancouver, Canada. Association for Computational Linguistics.
  • Chakrabarty et al. (2022) Abhisek Chakrabarty, Raj Dabre, Chenchen Ding, Hideki Tanaka, Masao Utiyama, and Eiichiro Sumita. 2022. FeatureBART: Feature based sequence-to-sequence pre-training for low-resource NMT. In Proceedings of the 29th International Conference on Computational Linguistics, pages 5014–5020, Gyeongju, Republic of Korea. International Committee on Computational Linguistics.
  • Chandra et al. (1981) Ashok K. Chandra, Dexter C. Kozen, and Larry J. Stockmeyer. 1981. Alternation. Journal of the Association for Computing Machinery, 28(1):114–133.
  • Chaudhary et al. (2018) Aditi Chaudhary, Chunting Zhou, Lori Levin, Graham Neubig, David R. Mortensen, and Jaime Carbonell. 2018. Adapting word embeddings to new languages with morphological and phonological subword representations. In Proceedings of the 2018 Conference on Empirical Methods in Natural Language Processing, pages 3285–3295, Brussels, Belgium. Association for Computational Linguistics.
  • Chercheur (1994) J.L. Chercheur. 1994. Case-Based Reasoning, 2nd edition. Morgan Kaufman Publishers, San Mateo, CA.
  • Chomsky (1973) N. Chomsky. 1973. Conditions on transformations. In A festschrift for Morris Halle, New York. Holt, Rinehart & Winston.
  • Chomsky (1993) Noam Chomsky. 1993. Lectures on government and binding: The Pisa lectures. 9. Walter de Gruyter.
  • Comrie (1989) Bernard Comrie. 1989. Language universals and linguistic typology: Syntax and morphology. University of Chicago press.
  • Conneau et al. (2018) Alexis Conneau, Ruty Rinott, Guillaume Lample, Adina Williams, Samuel R. Bowman, Holger Schwenk, and Veselin Stoyanov. 2018. Xnli: Evaluating cross-lingual sentence representations. In Proceedings of the 2018 Conference on Empirical Methods in Natural Language Processing. Association for Computational Linguistics.
  • Cooley and Tukey (1965) James W. Cooley and John W. Tukey. 1965. An algorithm for the machine calculation of complex Fourier series. Mathematics of Computation, 19(90):297–301.
  • Devlin et al. (2018) Jacob Devlin, Ming-Wei Chang, Kenton Lee, and Kristina Toutanova. 2018. Bert: Pre-training of deep bidirectional transformers for language understanding. arXiv preprint arXiv:1810.04805.
  • Dolan and Brockett (2005) William B. Dolan and Chris Brockett. 2005. Automatically constructing a corpus of sentential paraphrases. In Proceedings of the Third International Workshop on Paraphrasing (IWP2005).
  • Driess et al. (2023) Danny Driess, Fei Xia, Mehdi SM Sajjadi, Corey Lynch, Aakanksha Chowdhery, Brian Ichter, Ayzaan Wahid, Jonathan Tompson, Quan Vuong, Tianhe Yu, et al. 2023. Palm-e: An embodied multimodal language model. arXiv preprint arXiv:2303.03378.
  • Eco (1990) Umberto Eco. 1990. The Limits of Interpretation. Indian University Press.
  • Edmiston (2020) Daniel Edmiston. 2020. A systematic analysis of morphological content in bert models for multiple languages. arXiv preprint arXiv:2004.03032.
  • Ethayarajh (2019) Kawin Ethayarajh. 2019. How contextual are contextualized word representations? Comparing the geometry of BERT, ELMo, and GPT-2 embeddings. In Proceedings of the 2019 Conference on Empirical Methods in Natural Language Processing and the 9th International Joint Conference on Natural Language Processing (EMNLP-IJCNLP), pages 55–65, Hong Kong, China. Association for Computational Linguistics.
  • Gerz et al. (2018) Daniela Gerz, Ivan Vulić, Edoardo Maria Ponti, Roi Reichart, and Anna Korhonen. 2018. On the relation between linguistic typology and (limitations of) multilingual language modeling. In Proceedings of the 2018 Conference on Empirical Methods in Natural Language Processing, pages 316–327, Brussels, Belgium. Association for Computational Linguistics.
  • Glavaš and Vulić (2021) Goran Glavaš and Ivan Vulić. 2021. Is supervised syntactic parsing beneficial for language understanding tasks? an empirical investigation. In Proceedings of the 16th Conference of the European Chapter of the Association for Computational Linguistics: Main Volume, pages 3090–3104, Online. Association for Computational Linguistics.
  • Goldberg (2019) Yoav Goldberg. 2019. Assessing bert’s syntactic abilities. arXiv preprint arXiv:1901.05287.
  • Gusfield (1997) Dan Gusfield. 1997. Algorithms on Strings, Trees and Sequences. Cambridge University Press, Cambridge, UK.
  • Haspelmath (1999) Martin Haspelmath. 1999. Optimality and diachronic adaptation. Zeitschrift für Sprachwissenschaft, 18(2):180–205.
  • Haspelmath and Sims (2013) Martin Haspelmath and Andrea Sims. 2013. Understanding morphology. Routledge.
  • Hoel (1971a) Paul Gerhard Hoel. 1971a. Elementary Statistics, 3rd edition. Wiley series in probability and mathematical statistics. Wiley, New York, Chichester. ISBN 0 471 40300.
  • Hoel (1971b) Paul Gerhard Hoel. 1971b. Elementary Statistics, 3rd edition, Wiley series in probability and mathematical statistics, pages 19–33. Wiley, New York, Chichester. ISBN 0 471 40300.
  • Hofmann et al. (2020) Valentin Hofmann, Janet Pierrehumbert, and Hinrich Schütze. 2020. DagoBERT: Generating derivational morphology with a pretrained language model. In Proceedings of the 2020 Conference on Empirical Methods in Natural Language Processing (EMNLP), pages 3848–3861, Online. Association for Computational Linguistics.
  • Hofmann et al. (2021a) Valentin Hofmann, Janet Pierrehumbert, and Hinrich Schütze. 2021a. Dynamic contextualized word embeddings. In Proceedings of the 59th Annual Meeting of the Association for Computational Linguistics and the 11th International Joint Conference on Natural Language Processing (Volume 1: Long Papers), pages 6970–6984, Online. Association for Computational Linguistics.
  • Hofmann et al. (2021b) Valentin Hofmann, Janet Pierrehumbert, and Hinrich Schütze. 2021b. Superbizarre is not superb: Derivational morphology improves BERT’s interpretation of complex words. In Proceedings of the 59th Annual Meeting of the Association for Computational Linguistics and the 11th International Joint Conference on Natural Language Processing (Volume 1: Long Papers), pages 3594–3608, Online. Association for Computational Linguistics.
  • Hu et al. (2020) Junjie Hu, Sebastian Ruder, Aditya Siddhant, Graham Neubig, Orhan Firat, and Melvin Johnson. 2020. Xtreme: A massively multilingual multi-task benchmark for evaluating cross-lingual generalization. CoRR, abs/2003.11080.
  • Jawahar et al. (2019) Ganesh Jawahar, Benoît Sagot, and Djamé Seddah. 2019. What does BERT learn about the structure of language? In Proceedings of the 57th Annual Meeting of the Association for Computational Linguistics, pages 3651–3657, Florence, Italy. Association for Computational Linguistics.
  • Jespersen (1922) Otto Jespersen. 1922. Language: Its Nature, Development, and Origin. Allen and Unwin.
  • Kettunen (2014) Kimmo Kettunen. 2014. Can type-token ratio be used to show morphological complexity of languages? Journal of Quantitative Linguistics, 21(3):223–245.
  • Lan et al. (2019) Zhenzhong Lan, Mingda Chen, Sebastian Goodman, Kevin Gimpel, Piyush Sharma, and Radu Soricut. 2019. Albert: A lite bert for self-supervised learning of language representations. ArXiv, abs/1909.11942.
  • Lewkowycz et al. (2022) Aitor Lewkowycz, Anders Andreassen, David Dohan, Ethan Dyer, Henryk Michalewski, Vinay Ramasesh, Ambrose Slone, Cem Anil, Imanol Schlag, Theo Gutman-Solo, et al. 2022. Solving quantitative reasoning problems with language models. arXiv preprint arXiv:2206.14858.
  • Li et al. (2020) Xiaoya Li, Xiaofei Sun, Yuxian Meng, Junjun Liang, Fei Wu, and Jiwei Li. 2020. Dice loss for data-imbalanced NLP tasks. In Proceedings of the 58th Annual Meeting of the Association for Computational Linguistics, pages 465–476, Online. Association for Computational Linguistics.
  • Liu et al. (2019a) Yinhan Liu, Myle Ott, Naman Goyal, Jingfei Du, Mandar Joshi, Danqi Chen, Omer Levy, Mike Lewis, Luke Zettlemoyer, and Veselin Stoyanov. 2019a. Roberta: A robustly optimized BERT pretraining approach. CoRR, abs/1907.11692.
  • Liu et al. (2019b) Yinhan Liu, Myle Ott, Naman Goyal, Jingfei Du, Mandar Joshi, Danqi Chen, Omer Levy, Mike Lewis, Luke Zettlemoyer, and Veselin Stoyanov. 2019b. Roberta: A robustly optimized bert pretraining approach. arXiv preprint arXiv:1907.11692.
  • Martin et al. (2020) Louis Martin, Benjamin Muller, Pedro Javier Ortiz Suárez, Yoann Dupont, Laurent Romary, Éric de la Clergerie, Djamé Seddah, and Benoît Sagot. 2020. CamemBERT: a tasty French language model. In Proceedings of the 58th Annual Meeting of the Association for Computational Linguistics, pages 7203–7219, Online. Association for Computational Linguistics.
  • Mikolov et al. (2013) Tomas Mikolov, Ilya Sutskever, Kai Chen, Greg S Corrado, and Jeff Dean. 2013. Distributed representations of words and phrases and their compositionality. Advances in neural information processing systems, 26.
  • Müller (2002) Gereon Müller. 2002. Free word order, morphological case, and sympathy theory. Resolving Conflicts in Grammars: Optimality Theory in Syntax, Morphology, and Phonology. Buske, Hamburg, pages 9–48.
  • Nichols (1986) Johanna Nichols. 1986. Head-marking and dependent-marking grammar. Language, pages 56–119.
  • Nzeyimana and Niyongabo Rubungo (2022) Antoine Nzeyimana and Andre Niyongabo Rubungo. 2022. KinyaBERT: a morphology-aware Kinyarwanda language model. In Proceedings of the 60th Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers), pages 5347–5363, Dublin, Ireland. Association for Computational Linguistics.
  • Nzeyimana and Rubungo (2022) Antoine Nzeyimana and Andre Niyongabo Rubungo. 2022. Kinyabert: a morphology-aware kinyarwanda language model. arXiv preprint arXiv:2203.08459.
  • Otmakhova et al. (2022) Yulia Otmakhova, Karin Verspoor, and Jey Han Lau. 2022. Cross-linguistic comparison of linguistic feature encoding in BERT models for typologically different languages. In Proceedings of the 4th Workshop on Research in Computational Linguistic Typology and Multilingual NLP, pages 27–35, Seattle, Washington. Association for Computational Linguistics.
  • Park et al. (2021) Hyunji Hayley Park, Katherine J Zhang, Coleman Haley, Kenneth Steimel, Han Liu, and Lane Schwartz. 2021. Morphology matters: a multilingual language modeling analysis. Transactions of the Association for Computational Linguistics, 9:261–276.
  • Pennington et al. (2014) Jeffrey Pennington, Richard Socher, and Christopher D Manning. 2014. Glove: Global vectors for word representation. In Proceedings of the 2014 conference on empirical methods in natural language processing (EMNLP), pages 1532–1543.
  • Peters et al. (2018) Matthew E. Peters, Mark Neumann, Mohit Iyyer, Matt Gardner, Christopher Clark, Kenton Lee, and Luke Zettlemoyer. 2018. Deep contextualized word representations. In Proceedings of the 2018 Conference of the North American Chapter of the Association for Computational Linguistics: Human Language Technologies, Volume 1 (Long Papers), New Orleans, Louisiana. Association for Computational Linguistics.
  • Radford et al. (2019) Alec Radford, Jeffrey Wu, Rewon Child, David Luan, Dario Amodei, Ilya Sutskever, et al. 2019. Language models are unsupervised multitask learners. OpenAI blog, 1(8):9.
  • Raffel et al. (2020a) Colin Raffel, Noam Shazeer, Adam Roberts, Katherine Lee, Sharan Narang, Michael Matena, Yanqi Zhou, Wei Li, and Peter J. Liu. 2020a. Exploring the limits of transfer learning with a unified text-to-text transformer. J. Mach. Learn. Res., 21(1).
  • Raffel et al. (2020b) Colin Raffel, Noam Shazeer, Adam Roberts, Katherine Lee, Sharan Narang, Michael Matena, Yanqi Zhou, Wei Li, and Peter J Liu. 2020b. Exploring the limits of transfer learning with a unified text-to-text transformer. The Journal of Machine Learning Research, 21(1):5485–5551.
  • Rajpurkar et al. (2016) Pranav Rajpurkar, Jian Zhang, Konstantin Lopyrev, and Percy Liang. 2016. SQuAD: 100,000+ questions for machine comprehension of text. In Proceedings of the 2016 Conference on Empirical Methods in Natural Language Processing, pages 2383–2392, Austin, Texas. Association for Computational Linguistics.
  • Rasooli and Tetreault (2015) Mohammad Sadegh Rasooli and Joel R. Tetreault. 2015. Yara parser: A fast and accurate dependency parser. Computing Research Repository, arXiv:1503.06733. Version 2.
  • Sadeniemi et al. (2008) Markus Sadeniemi, Kimmo Kettunen, Tiina Lindh-Knuutila, and Timo Honkela. 2008. Complexity of european union languages: A comparative approach. Journal of Quantitative Linguistics, 15(2):185–211.
  • Sapir (1921) Edward Sapir. 1921. Language, an Introduction to the Study of Speech, by Edward Sapir. New York:: Harcourt, Brace.
  • Sharma et al. (2019) Lakshay Sharma, Laura Graesser, Nikita Nangia, and Utku Evci. 2019. Natural language understanding with the quora question pairs dataset. ArXiv, abs/1907.01041.
  • Shaw et al. (2018) Peter Shaw, Jakob Uszkoreit, and Ashish Vaswani. 2018. Self-attention with relative position representations. arXiv preprint arXiv:1803.02155.
  • Singer et al. (1954–58) Charles Joseph Singer, E. J. Holmyard, and A. R. Hall, editors. 1954–58. A history of technology. Oxford University Press, London. 5 vol.
  • Socher et al. (2013) Richard Socher, Alex Perelygin, Jean Wu, Jason Chuang, Christopher D. Manning, Andrew Ng, and Christopher Potts. 2013. Recursive deep models for semantic compositionality over a sentiment treebank. In Proceedings of the 2013 Conference on Empirical Methods in Natural Language Processing, pages 1631–1642, Seattle, Washington, USA. Association for Computational Linguistics.
  • Strötgen and Gertz (2012) Jannik Strötgen and Michael Gertz. 2012. Temporal tagging on different domains: Challenges, strategies, and gold standards. In Proceedings of the Eight International Conference on Language Resources and Evaluation (LREC’12), pages 3746–3753, Istanbul, Turkey. European Language Resource Association (ELRA).
  • Su et al. (2021) Jianlin Su, Yu Lu, Shengfeng Pan, Ahmed Murtadha, Bo Wen, and Yunfeng Liu. 2021. Roformer: Enhanced transformer with rotary position embedding. arXiv preprint arXiv:2104.09864.
  • Superman et al. (2000) S. Superman, B. Batman, C. Catwoman, and S. Spiderman. 2000. Superheroes experiences with books, 20th edition. The Phantom Editors Associates, Gotham City.
  • Tenney et al. (2019) Ian Tenney, Dipanjan Das, and Ellie Pavlick. 2019. BERT rediscovers the classical NLP pipeline. In Proceedings of the 57th Annual Meeting of the Association for Computational Linguistics, pages 4593–4601, Florence, Italy. Association for Computational Linguistics.
  • Vaswani et al. (2017) Ashish Vaswani, Noam Shazeer, Niki Parmar, Jakob Uszkoreit, Llion Jones, Aidan N Gomez, Łukasz Kaiser, and Illia Polosukhin. 2017. Attention is all you need. Advances in neural information processing systems, 30.
  • Wang et al. (2018) Alex Wang, Amanpreet Singh, Julian Michael, Felix Hill, Omer Levy, and Samuel R Bowman. 2018. Glue: A multi-task benchmark and analysis platform for natural language understanding. arXiv preprint arXiv:1804.07461.
  • Warstadt et al. (2018) Alex Warstadt, Amanpreet Singh, and Samuel R Bowman. 2018. Neural network acceptability judgments. arXiv preprint arXiv:1805.12471.
  • Williams et al. (2018) Adina Williams, Nikita Nangia, and Samuel Bowman. 2018. A broad-coverage challenge corpus for sentence understanding through inference. In Proceedings of the 2018 Conference of the North American Chapter of the Association for Computational Linguistics: Human Language Technologies, Volume 1 (Long Papers), pages 1112–1122. Association for Computational Linguistics.
  • Wolf et al. (2019) Thomas Wolf, Lysandre Debut, Victor Sanh, Julien Chaumond, Clement Delangue, Anthony Moi, Pierric Cistac, Tim Rault, Rémi Louf, Morgan Funtowicz, et al. 2019. Huggingface’s transformers: State-of-the-art natural language processing. arXiv preprint arXiv:1910.03771.
  • Xue et al. (2020) Linting Xue, Noah Constant, Adam Roberts, Mihir Kale, Rami Al-Rfou, Aditya Siddhant, Aditya Barua, and Colin Raffel. 2020. mt5: A massively multilingual pre-trained text-to-text transformer. In North American Chapter of the Association for Computational Linguistics.
  • Yang et al. (2019) Yinfei Yang, Yuan Zhang, Chris Tar, and Jason Baldridge. 2019. PAWS-X: A cross-lingual adversarial dataset for paraphrase identification. In Proceedings of the 2019 Conference on Empirical Methods in Natural Language Processing and the 9th International Joint Conference on Natural Language Processing (EMNLP-IJCNLP), pages 3687–3692, Hong Kong, China. Association for Computational Linguistics.
  • Zeman et al. (2023) Daniel Zeman et al. 2023. Universal dependencies 2.12. LINDAT/CLARIAH-CZ digital library at the Institute of Formal and Applied Linguistics (ÚFAL), Faculty of Mathematics and Physics, Charles University.